# DN-4DGS: Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering

### v1

@@ 5.3 Ablation Studies
TiNeuVox [45] 21.61 0.823 0.277
HyperNeRF [2] 23.45 0.849 0.199
NeRF-DS [47] 23.60 0.849 0.182
3D-GS [12] 20.29 0.782 0.292
D3DGS [13] 23.92 0.847 0.184
@@ 5.4 Limitations and Future Work
Although two-stage deformation can alter the coordinate distribution of canonical 3D gaussians and reduce the noise introduced into the deformation field, the lack of simultaneous supervision for both stages [52; 53; 41] poses a challenge. Consequently, during the second stage, due to the lack of supervision in the first-stage deformation, the direction of coordinate deformation becomes uncontrollable to some extent. This, in turn, affects the spatial feature aggregation in the second stage. To address this issue, future work should explore the direction of simultaneous supervision.
@@ A.2 More Model Details
Feature Encoding. As illustrated in Section 4.3.1, we can utilize the MLPs from D3DGS [13] or the HexPlanes from 4DGaussian [11] as Feature Encoding. Specifically, for both PlenopticVideo [1] and HyperNeRF [2], we use the HexPlanes to encode per-gaussian’s feature. The complete details can be referred to 4DGaussian’s main text. For NeRF-DS [47], we utilize the MLPs for feature encoding. The complete details can be referred to D3DGS’s main text.
@@ A.5 Detailed Results
NeRFPlayer [35] 31.83 0.928 32.06 0.930 32.31 0.940
HexPlane [15] 32.71 0.985 31.86 0.983 32.09 0.986
KPlanes [34] 31.82 0.966 32.60 0.966 32.52 0.974
MixVoxels [31] 31.30 0.965 31.65 0.965 31.43 0.971
4DGaussian [11] 32.90 0.957 32.46 0.949 32.49 0.957
@@
NeRFPlayer [35] 27.36 0.867 26.14 0.849 32.05 0.938
HexPlane [15] 31.92 0.988 29.26 0.980 - -
KPlanes [34] 32.39 0.970 30.44 0.953 29.99 0.953
MixVoxels [31] 31.21 0.970 29.92 0.945 29.36 0.946
4DGaussian [11] 32.51 0.954 29.20 0.917 27.34 0.905
@@
Nerfies [19] 20.6 0.83 26.7 0.94 19.2 0.56 22.4 0.87
HyperNeRF [2] 20.0 0.59 26.9 0.94 19.3 0.59 23.3 0.90
TiNeuVox-B [45] 22.8 0.84 28.3 0.95 21.5 0.69 24.4 0.87
FFDNeRF [54] 22.8 0.84 28.0 0.94 21.9 0.71 24.3 0.86
3D-GS [12] 18.3 0.60 19.7 0.70 20.6 0.63 20.4 0.80
4DGaussian [11$\ddagger$ 20.9 0.75 24.1 0.82 20.0 0.53 22.2 0.74
@@
3D-GS [12] 23.16 16.14 21.01 22.89 21.71 22.69 18.42
TiNeuVox [45] 21.49 20.58 23.08 24.47 19.71 21.26 20.66
HyperNeRF [2] 25.43 18.93 23.06 26.15 24.59 25.58 20.41
NeRF-DS [47] 25.78 20.54 23.19 25.72 24.91 25.13 19.96
D3DGS [13] 25.72 20.40 25.24 25.70 24.35 26.35 19.70
@@ A.7 Assets Availability
The datasets that support the findings of this study are available in the following repositories: HyperNeRF [2] at https://github.com/google/hypernerf/releases/tag/v0.1 under Apache-2.0 license. NeRF-DS [47] at https://github.com/JokerYan/NeRF-DS/releases/tag/v0.1-pre-release under Apache-2.0 license. PlenopticVideo [1] at https://github.com/facebookresearch/Neural_3D_Video?tab=License-1-ov-file under CC BY-NC 4.0 license. The code of our baseline [11; 13] is available at https://github.com/ingra14m/Deformable-3D-Gaussians under MIT license and https://github.com/hustvl/4DGaussians under Gaussian-Splatting license.

### v2

@@ 5.3 Ablation Studies
TiNeuVox tineuvox 21.61 0.823 0.277
HyperNeRF park2021hypernerf 23.45 0.849 0.199
NeRF-DS yan2023nerf 23.60 0.849 0.182
3D-GS 3dgs 20.29 0.782 0.292
D3DGS yang2023deformable 23.92 0.847 0.184
@@ 5.4 Limitations and Future Work
Although two-stage deformation can alter the coordinate distribution of canonical 3D gaussians and reduce the noise introduced into the deformation field, the lack of simultaneous supervision for both stages carion2020end ; cheng2022masked ; lu2023query poses a challenge. Consequently, during the second stage, due to the lack of supervision in the first-stage deformation, the direction of coordinate deformation becomes uncontrollable to some extent. This, in turn, affects the spatial feature aggregation in the second stage. To address this issue, future work should explore the direction of simultaneous supervision.
@@ A.2 More Model Details
Feature Encoding. As illustrated in Section 4.3.1, we can utilize the MLPs from D3DGS yang2023deformable or the HexPlanes from 4DGaussian wu20234d as Feature Encoding. Specifically, for both PlenopticVideo li2022neural and HyperNeRF park2021hypernerf , we use the HexPlanes to encode per-gaussian’s feature. The complete details can be referred to 4DGaussian’s main text. For NeRF-DS yan2023nerf , we utilize the MLPs for feature encoding. The complete details can be referred to D3DGS’s main text.
@@ A.5 Detailed Results
NeRFPlayer song2023nerfplayer 31.83 0.928 32.06 0.930 32.31 0.940
HexPlane hexplane 32.71 0.985 31.86 0.983 32.09 0.986
KPlanes kplanes 31.82 0.966 32.60 0.966 32.52 0.974
MixVoxels wang2023mixedvoxels 31.30 0.965 31.65 0.965 31.43 0.971
4DGaussian wu20234d 32.90 0.957 32.46 0.949 32.49 0.957
@@
NeRFPlayer song2023nerfplayer 27.36 0.867 26.14 0.849 32.05 0.938
HexPlane hexplane 31.92 0.988 29.26 0.980 - -
KPlanes kplanes 32.39 0.970 30.44 0.953 29.99 0.953
MixVoxels wang2023mixedvoxels 31.21 0.970 29.92 0.945 29.36 0.946
4DGaussian wu20234d 32.51 0.954 29.20 0.917 27.34 0.905
@@
Nerfies park2021nerfies 20.6 0.83 26.7 0.94 19.2 0.56 22.4 0.87
HyperNeRF park2021hypernerf 20.0 0.59 26.9 0.94 19.3 0.59 23.3 0.90
TiNeuVox-B tineuvox 22.8 0.84 28.3 0.95 21.5 0.69 24.4 0.87
FFDNeRF guo2023forwardflowfornvs 22.8 0.84 28.0 0.94 21.9 0.71 24.3 0.86
3D-GS 3dgs 18.3 0.60 19.7 0.70 20.6 0.63 20.4 0.80
4DGaussian wu20234d $\ddagger$ 20.9 0.75 24.1 0.82 20.0 0.53 22.2 0.74
@@
3D-GS 3dgs 23.16 16.14 21.01 22.89 21.71 22.69 18.42
TiNeuVox tineuvox 21.49 20.58 23.08 24.47 19.71 21.26 20.66
HyperNeRF park2021hypernerf 25.43 18.93 23.06 26.15 24.59 25.58 20.41
NeRF-DS yan2023nerf 25.78 20.54 23.19 25.72 24.91 25.13 19.96
D3DGS yang2023deformable 25.72 20.40 25.24 25.70 24.35 26.35 19.70
@@ A.7 Assets Availability
The datasets that support the findings of this study are available in the following repositories: HyperNeRF park2021hypernerf at https://github.com/google/hypernerf/releases/tag/v0.1 under Apache-2.0 license. NeRF-DS yan2023nerf at https://github.com/JokerYan/NeRF-DS/releases/tag/v0.1-pre-release under Apache-2.0 license. PlenopticVideo li2022neural at https://github.com/facebookresearch/Neural_3D_Video?tab=License-1-ov-file under CC BY-NC 4.0 license. The code of our baseline wu20234d ; yang2023deformable is available at https://github.com/ingra14m/Deformable-3D-Gaussians under MIT license and https://github.com/hustvl/4DGaussians under Gaussian-Splatting license.